A particle moves so that its position vector at time t is given by r =e−t(costi +sintj). Show that at any time t, (a) its velocity v is inclined to the vector r at a constant angle 3 (b) its acceleration vector is at right angles to the vector r.


1
Expert's answer
2021-03-31T01:52:22-0400
"\text{r}=(e^{-t}\cos t)\text{i}+(e^{-t}\sin t)\text{j}"

"\text{v}=(-e^{-t}\cos t-e^{-t}\sin t)\text{i}"

"+(-e^{-t}\sin t+e^{-t}\cos t)\text{j}"


"\text{a}=(2e^{-t}\sin t)\text{i}+(-2e^{-t}\cos t)\text{j}"

a)

"\text{r}\cdot \text{v}=(e^{-t}\cos t)(-e^{-t}\cos t-e^{-t}\sin t)"

"+(e^{-t}\sin t)(-e^{-t}\sin t+e^{-t}\cos t)"

"=-e^{-2t}\cos^2t-e^{-2t}\cos t\sin t"

"-e^{-2t}\sin^2t+e^{-2t}\cos t\sin t=-e^{-2t}"


"|\text{r}|=\sqrt{(e^{-t}\cos t)^2+(e^{-t}\sin t)^2 }=e^{-t}"

"(-e^{-t}\cos t-e^{-t}\sin t)^2+(-e^{-t}\sin t+e^{-t}\cos t)^2"

"=e^{-2t}\cos^2t+2e^{-2t}\cos t\sin t+e^{-2t}\sin^2t"

"e^{-2t}\cos^2t-2e^{-2t}\cos t\sin t+e^{-2t}\sin^2t=2e^{-2t}"


"|\text{v}|=\sqrt{2e^{-2t} }=\sqrt{2}e^{-t}"

"\cos\angle(\text{r}, \text{v})=\dfrac{\text{r}\cdot \text{v}}{|\text{r}||\text{v}|}"

"=\dfrac{-e^{-2t}}{e^{-t}\sqrt{2}e^{-t}}=-\dfrac{\sqrt{2}}{2}"

"\angle(\text{r}, \text{v})=\dfrac{3\pi}{4}=\text{const}"



b)


"\text{r}\cdot \text{a}=(e^{-t}\cos t)(2e^{-t}\sin t)"

"+(e^{-t}\sin t)(-2e^{-t}\cos t)=0"

Hence acceleration vector "\text{a}" is at right angles to the vector "\text{r}".